Circular No. 7926

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

1E 2259+586 V. M. Kaspi, McGill University and Massachusetts Institute of Tech-

nology; F. P. Gavriil, McGill University; and P. M. Woods, National Space Science and Technology Center, report on further analysis of RXTE/PCA data from 1E 2259+586 (cf. IAUC 7924), obtained during its outburst on June 18: “The pulsed flux in the range 2–10 keV declined monotonically by a factor of 4 over the 15000 s in which the bursts were detected. The burst rate decreased similarly. The pulse profile during the 15000 s was significantly different than the long-term average profile, an effect not seen in 5.8 yr of prior RXTE monitoring. These two observations independently demonstrate that 1E 2259+586 is the origin of the bursts, solidifying the common nature of anomalous x-ray pulsars and soft-γ-ray repeaters. Bursting in anomalous x-ray pulsars was predicted in the magnetar model (Thompson and Duncan 1996, Ap.J. 473, 322).”

Kaspi also reports (together with J. Jensen, F. Rigaut, and A. Hatakeyama, Gemini Observatory; and P. M. Woods) that near-infrared observations of 1E 2259+586 were obtained on June 21.614 UT with the Gemini North 8-m reflector (+ NIRI + K_{s} filter). In a 30-min exposure (seeing 0′′.7), the reported possible infrared counterpart to the pulsar (Hulleman et al. 2000, A.Ap. 358, 605) was detected with preliminary magnitude K_{s} = 20.36 ± 0.15 (1.3 mag brighter than the quiescent value), indicating that, ∼ 3 days following the bursts detected with RXTE/PCA (IAUC 7924), the source was a factor of 3.4 brighter (5.4σ confidence limit), thus confirming the infrared identification. Further infrared/optical monitoring is advised.

THE EDGAR WILSON AWARD The Smithsonian Astrophysical Observatory announces that the 2002

Edgar Wilson Award for the discovery of comets (cf. IAUC 7652), is being divided among the following seven individuals: Vance Avery Petriew, Regina, SK, Canada, for P/2001 Q2; Kaoru Ikeya, Mori, Shuchi, Shizuoka, Japan, and Daqing Zhang, Kaifeng, Henan province, China, for C/2002 C1; Douglas Snyder, Palominas, AZ, U.S.A., and Shigeki Murakami, Matsunoyama, Niigata, Japan, for C/2002 E2; Syogo Utsunomiya, Minami-Oguni, Aso, Kumamoto, Japan, for C/2002 F1; William Kwong Yu Yeung, Benson, AZ, U.S.A., for P/2002 BV.

2002 June 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green